OpenStudy (anonymous):

Two point charges are seperated by a distance R in air repel each other by a force F.One of the charges is doubled and the other is made one - fourth.For what new seperation between them will the force remain same?

OpenStudy (anonymous):

\[1/\sqrt{2}\]R whre R is the distance between two charges in case one

OpenStudy (anonymous):

\[F=KQ _{1} Q _{2}/R ^{2}\] \[F´=K2Q _{1} Q _{2}/4R ´^{2}\] FOR F=F´ THEN \[KQ _{1} Q _{2}/R ^{2}=K2Q _{1} Q _{2}/4R´ ^{2}\] \[1/R ^{2}=2/4R´ ^{2}\] \[R´ ^{2}=R ^{2}/2\] \[R´=R \sqrt{2}/2\]